OpenStudy (anonymous):

which of these cannot represent the lengths of the sides of a right triangle A. 3ft, 4ft, 5 ft B. 6in , 8in , 10in C. 16 cm, 63 cm, 65 cm, D. 8m , 9m , 10m

jhonyy9 (jhonyy9):

so what you know about the lengths of a right triangle in function of pythagora"s theorem ?

jhonyy9 (jhonyy9):

the square of hypothenuse need being equal with the sum of ... do you know this ?

OpenStudy (anonymous):

yup

jhonyy9 (jhonyy9):

so than try these choises A. 3,4 and 5 this mean that 9 +16 =25 so this is right hope you can continue and will get the right answer

jhonyy9 (jhonyy9):

so B. 36 +64 =100 is right too yes ?

jhonyy9 (jhonyy9):

ok so what you think abour D. 64 +81 =100 so this is right ?

jhonyy9 (jhonyy9):

not is sure

jhonyy9 (jhonyy9):

ok now you need calculi the choises C. is right or not ?

jhonyy9 (jhonyy9):

courage please or you dont like cooperat nothing ?

jhonyy9 (jhonyy9):

16,63 and 65 what you get ?

OpenStudy (anonymous):

i add it right

jhonyy9 (jhonyy9):

not 16 squared = ? 63 squared = ? 65 squared = ? and you need to see than the sum of 16 squared and 63 squared will be equale with 65 squared ok ?

jhonyy9 (jhonyy9):

or you dont know what mean the pythagora"s theorem ?

OpenStudy (anonymous):

well the square root of 16 is 4 63=7.937253933193772 65=8.06225774829855

jhonyy9 (jhonyy9):

not the squareroot the square like 16^2 = ?

OpenStudy (anonymous):

lol sorry

jhonyy9 (jhonyy9):

try it now ,courage

jhonyy9 (jhonyy9):

16^2 =256 continue

OpenStudy (anonymous):

and 63 is 25047

jhonyy9 (jhonyy9):

63^2 = 3969 and 65^2 = sorry but i dont know how you have got this

jhonyy9 (jhonyy9):

so 65^2 = 4225

OpenStudy (anonymous):

sorry it cuz i pushed the wrong numbers

jhonyy9 (jhonyy9):

so than 256 + 3969 = 4225 ????

jhonyy9 (jhonyy9):

so yes is right

jhonyy9 (jhonyy9):

than you see that the choise D not is right sure hope so much that you understand it now sure yes ?

OpenStudy (anonymous):

yup thanks a lot and thanks for the patience

jhonyy9 (jhonyy9):

was my pleasure good luck bye